OpenStudy (anonymous):

explain in complete sentences how you would use the SUBSTITUTION METHOD to solve the following system of equations. 5x - y = -1 7x - 3y = 13

OpenStudy (anonymous):

can someone please answer this problem im very confused

OpenStudy (anonymous):

yea sure :)

OpenStudy (anonymous):

ok thank you

OpenStudy (anonymous):

-y=-1-5x y=1+5 now substitute y in 2nd equation

OpenStudy (anonymous):

7x-3(1+5x)=13

OpenStudy (anonymous):

is that last subsitution to check the answere?

OpenStudy (anonymous):

In the Savannah, a sinkhole opened catching some ostrich and cheetah. There was a total of 13 animals to be rescued. If 44 INDIVIDUAL feet were pulled from the muck, how many ostrich and cheetah were rescued?

OpenStudy (anonymous):

last question i need help with

OpenStudy (anonymous):

seems easy but i can't seem to figure it out

OpenStudy (anonymous):

7x-3-15x=13 -8x=13+3 x=16/-8 x=-2 now,substitute x value in one of the equation 5(-2)-y=-1 -10-y=-1 -y=-1+10 -y=9 y=-9

OpenStudy (anonymous):

its for finding the value of y

OpenStudy (anonymous):

ok.

jhonyy9 (jhonyy9):

from first equation you get y=5x+1 7x-3y=13 --- 7x-3(5x+1)=13 --- 7x-15x-3=13 --- -8x = 16 --- x=16/(-8) x= -2 --- than y=5(-2) +1 --- y= -10+1 --- y= -9 x= -2 y= -9